# Frobenius intertwiners for $q$-difference equations

### v1

@@ 9. Concluding remarks and future directions
It is known that the Nakajima quiver varieties have very rich enumerative geometry [Oko17]. Partition functions of quasimaps to the Nakajima quiver varieties with various boundary conditions satisfy remarkable systems of $q$-difference equations, known as the quantum difference equations [OS22]. In the present paper we discuss the existence of the Frobenuis structure for the quantum difference equations associated with simplest Nakajima variety $X=T^{*}\mathbb{P}^{n-1}$. All essential ingredients in our construction, e.g., the existence of the rational shift operators in the equivariant parameter $\hbar$, exist for general Nakajima varieties too. For this reason, the $p$ adic quantum difference equations of the Nakajima varieties are equipped with the Frobenius structure. We plan to return to the case of general $X$ in a separate note. In the cohomological limit $q\to 1$ the quantum difference equations degenerate to the quantum differential equations for the Nakajima varieties [MO19], therefore the Frobenius structures exists for such differential equations also.
@@
The quantum difference/differential equations of the Nakajima varieties are connected deeply with representation theory of quantum groups. It is an interesting problem to find a representation theoretic description of the Frobenius intertwiners for these equations. In particular, the quantum difference equations for $q$ specialised at the roots of unity are controlled by the Bethe ansatz, see [AO17, PSZ, KPSZ]. It seems possible to describe the Frobenius intertwiners for these values of $q$ as transition matrices between various Bethe eigenvectors. We plan to investigate this question further in [KS].
9.3.
In the framework of enumerative geometry, one can define the Frobenius intertwiner as the partition function counting quasimaps from $\mathbb{P}^{1}$ to a Nakajima variety $X$ with two different relative conditions at $0\in\mathbb{P}^{1}$ and $\infty\in\mathbb{P}^{1}$. The corresponding moduli space of quasimaps is represented by Fig.1, we refer to [Oko17] and also [PSZ] where the pictorial notations are explained.
At $0\in\mathbb{P}^{1}$ we have the usual relative condition on quasimaps as defined in [Oko17]. We recall that the relative boundary condition at $0\in\mathbb{P}^{1}$ means that $\mathbb{P}^{1}$ is allowed to have a chain of rational “bubbles” with simple nodal intersections. This chain of bubbles is represented by blue color in Fig. 1.
At $\infty\in\mathbb{P}^{1}$ we also impose a relative condition but require, in addition, the relative quasimaps to be $\mathbb{Z}/p\mathbb{Z}$-invariant. More precisely, recall that the moduli space of relative quasimaps is defined as a stack quotient of quasimaps from a chain of $\mathbb{P}^{1}$’s by the automorphism group (×$(\mathbb{C}^{\times})^{l}$ where $l$ is the length of the chain, see Section 6.2 in [Oko17]. The moduli space of relative quasimaps at $\infty\in\mathbb{P}^{1}$ is constructed as follows: first we consider the space of quasimaps from a chain of rational curves of length $l$ which are invariant with respect to $\mathbb{Z}/p\mathbb{Z}$, which acts in each “bubble” (the component of the chain) via multiplication by $p$ th roots of unity. Then we define the $\mathbb{Z}/p\mathbb{Z}$-invariant quasimaps as the stack quotient of this moduli space by (×$(\mathbb{C}^{\times})^{l}$.
It is known, Theorem 8.1.16 in [Oko17], that the partition function of relative quasimaps from the chain of bubbles at $0\in\mathbb{P}^{1}$ is a fundamental solution matrix $\Psi({\boldsymbol{u}},\hbar,q,z)$ of the quantum difference equation associated with $X$, normalized so that
$\Psi({\boldsymbol{u}},\hbar,q,z)=1+O(z),$
Similarly, the partition function of $\mathbb{Z}/p\mathbb{Z}$ invariant quasimaps at $\infty\in\mathbb{P}^{1}$ can be shown to be equal to $\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$. Finally, the contribution of degree 00 - quasimaps which correspond to the “constant” quasimaps from the parameterized component $\mathbb{P}^{1}$ can be computed by $\mathbb{C}^{\times}_{q}$ equivariant localization similarly to how it was done for the shift operators in Section 8.2 of [Oko17], in particular see Lemma 8.2.12. The resulting contribution has the form
$\mathsf{U}(0)=\Gamma_{p,q}\Big{(}(q-\hbar)T^{1/2}X\Big{)},$
where $T^{1/2}X$ denotes the corresponding choice of the polarization bundle of $X$. For instance, in the case $X=T^{*}\mathbb{P}^{1}$ this formula specializes to (3.5). Combining all this together we obtain that the corresponding partition function is
$\mathsf{U}(z)=\Psi({\boldsymbol{u}},\hbar,q,z)\mathsf{U}(0)\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$
which is exactly the Frobenius intertwiner.
@@
We note that the idea of $\mathbb{Z}/p\mathbb{Z}$ invariant quasimap moduli is similar to the moduli space recently used in [HL24] to define the quantum Steenrod operators in quantum cohomology.
9.4.
Frobenius intertwiners link the theory of $p$ adic differential equations with the theory of exponential sums in number theory [K90]. Namely, interesting sums arise as the traces
(9.1) $\textrm{Tr}(\mathsf{U}(T_{z}))$
where $T_{z}\in\mathbb{Q}_{p}$ denotes the Teichmuller representative of an element $z\in\mathbb{F}_{p}$. These sums may be considered as the arithmetic counterpart of the integral representation for solutions of the corresponding differential equation. See for instance [Dwo74], where the Kloosterman sum is realized as the trace of Frobenius for the Bessel differential equation.
These type of sums for the $q$ case seem never considered. One interesting new feature of the $q$ difference equations for Nakajima varieties is the $3D$-mirror symmetry, which is the invariance of the equations with respect to transformation $z\longleftrightarrow{\boldsymbol{u}}$, where $z$ denote the Kähler parameters (parameters counting the degrees of the quasimaps) and ${\boldsymbol{u}}$ denote the equivariant parameters. It is interesting to investigate what part of this symmetry survives in the arithmetic context of the exponential sums (9.1).
@@ Acknowledgements
I thank K. Kedlaya, J. Hee Lee, A. Oblomkov, D. Vargas-Montoya, M. Vlasenko and especially A. Okounkov for very inspiring discussions. I thank A. Varchenko for introducing me to the subject. This note is the extended version of the talks [Sm23] and [Sm24], I thank the organizers for these opportunities. This work is supported by NSF grants DMS - 2054527 and DMS-2401380.

### v2

@@ 9. Concluding remarks and future directions
We expect that the Frobenius intertwiner can be defined enumeratively as a partition function counting the quasimaps from $\mathbb{P}^{1}$ to a Nakajima variety $X$. For this, one can consider the quasimaps with two different relative conditions at $0\in\mathbb{P}^{1}$ and $\infty\in\mathbb{P}^{1}$. The corresponding moduli space of quasimaps is represented by Fig.1, we refer to [Oko17] and also [PSZ] where the pictorial notations for various quasimap moduli spaces were introduced.
At $0\in\mathbb{P}^{1}$ we have the usual relative condition on quasimaps as defined in [Oko17]. We recall that the relative boundary condition at $0\in\mathbb{P}^{1}$ means that $\mathbb{P}^{1}$ is allowed to deform into a chain of rational “bubbles” with simple nodal intersections. This chain of bubbles is represented by the blue color in Fig. 1.
It is known, Theorem 8.1.16 in [Oko17], that the partition function of relative quasimaps from the chain of bubbles at $0\in\mathbb{P}^{1}$ is a fundamental solution matrix $\Psi(\hbar,{\boldsymbol{u}},q,z)$ of the quantum difference equation associated with $X$, normalized so that
$\Psi(\hbar,{\boldsymbol{u}},q,z)=1+O(z)\in K_{T}(X)^{\otimes 2}[[z]],$
At $\infty\in\mathbb{P}^{1}$ we also impose a relative condition but require, in addition, the relative quasimaps to be $\mathbb{Z}/p\mathbb{Z}$-invariant. More precisely, recall that the moduli space of relative quasimaps is defined as a stack quotient of quasimaps from a chain of $\mathbb{P}^{1}$’s by the automorphism group (×$(\mathbb{C}^{\times})^{l}$ where $l$ is the length of the chain, see Section 6.2 in [Oko17]. The moduli space of relative quasimaps at $\infty\in\mathbb{P}^{1}$ is constructed as follows: first we consider the space of quasimaps from a chain of rational curves of length $l$ which are invariant with respect to $\mathbb{Z}/p\mathbb{Z}$, which acts in each “bubble” (the component of the chain) via multiplication by $p$ th roots of unity. Then we define the $\mathbb{Z}/p\mathbb{Z}$-invariant quasimaps as the stack quotient of this moduli space by (×$(\mathbb{C}^{\times})^{l}$. This “$\mathbb{Z}/p\mathbb{Z}$-invariant chain of bubbles” is represented by the red color in Fig. 1.
By the $\mathbb{Z}/p\mathbb{Z}$-invariance, the degrees of the quasimaps allowed in this picture are all multiples of $p$. In particular, the corresponding partition function is a power series in $z^{p}$. For appropriately defined moduli space, the corresponding partition function arising at $\infty\in\mathbb{P}^{1}$ is equal to $\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$. The inverse arises due to transposition and inversion of character $q\to 1/q$ at $\infty\in\mathbb{P}^{1}$.
Finally, the contribution of degree 00 - quasimaps which correspond to the “constant” quasimaps from the parameterized component $\mathbb{P}^{1}$ can be computed by $\mathbb{C}^{\times}_{q}$ equivariant localization similarly to how it was done for the shift operators in Section 8.2 of [Oko17], in particular see Lemma 8.2.12. The resulting contribution has the form
$\dfrac{\Phi(ph,p\boldsymbol{a},px,q,z)}{\Phi(h,\boldsymbol{a},x,q^{p},z^{p})}\sim\Gamma_{p,q}\Big{(}(q^{p}-\hbar^{p})T^{1/2}X^{(p)}\Big{)},$
where $T^{1/2}X$ denotes the corresponding choice of the polarization bundle of $X$. The numerator of this fraction is the localization contribution at $0\in\mathbb{P}^{1}$ and the denominator is at $\infty\in\mathbb{P}^{1}$. Combining all this together we obtain that the corresponding partition function is
$\Psi({\boldsymbol{u}},\hbar,q,z)\Gamma_{p,q}\Big{(}(q^{p}-\hbar^{p})T^{1/2}X^{(p)}\Big{)}\Psi({\boldsymbol{u}},\hbar,q^{p},z^{p})^{-1}$
which is exactly the Frobenius intertwiner (5.2).
@@
The idea of $\mathbb{Z}/p\mathbb{Z}$ invariant quasimap moduli is similar to the one used in [HL24] to investigate the quantum Steenrod operators in quantum cohomology. It was shown in [HL24] that for a large class of varieties $X$, which includes the Nakajima varieties, the quantum Steenrod operations coincide with the $p$-curvature of the quantum connection.
The $q$-difference connections also have well defined $p$-curvature. Namely, the analog of $p$-curvature of $q$-difference connection (2.10) is:
$\mathbf{C}^{(p)}(\zeta,z)=\left.\Big{(}\mathbf{M}(zq^{p-1})\dots\mathbf{M}(zq)\mathbf{M}(z)\Big{)}\right|_{q=\zeta}$
where $\zeta$ is a non-trivial unit root of order $p$. In [KS] we showed that the $p$-curvature of the quantum connection of $X$ arises as the first non-trivial term in the $p$-adic expansion of $\mathbf{C}^{(p)}(\zeta,z)$.
Let us show that $p$-curvature $\mathbf{C}^{(p)}(\zeta,z)$ is similar to $\mathbf{M}(z^{p})$. Indeed, iterating the $q$-difference equation (2.10) $p$-times we obtain:
$\Psi(q,zq^{p})=\mathbf{M}(zq^{p-1})\dots\mathbf{M}(z)\Psi(z)$
also, rising all variables in (2.10) to their $p$-th powers we obtain:
$\Psi(q^{p},z^{p}q^{p})=\mathbf{M}(z^{p})\Psi(q^{p},z^{p})$
Dividing the first equation by the second, and specializing at $q=\zeta$ we arrive at:
$\mathsf{U}(\zeta,z)\mathbf{M}(z^{p})\mathsf{U}(\zeta,z)^{-1}=\mathbf{C}^{(p)}(\zeta,z).$
In particular, the spectrum of the $q$-deformed $p$-curvature $\mathbf{C}^{(p)}(\zeta,z)$ coincides with the spectrum of $\psi^{p}$-twisted quantum multiplication $\mathbf{M}(z^{p})$. This result is a $q$-version of the isospectrality recently proven in [EV24] in the differential case using different methods.
@@ Acknowledgements
I thank S. Bai, P. Etingof, V. Golyshev, K. Kedlaya, J. Hee Lee, A. Oblomkov, A. Petrov, D. Vargas-Montoya, M. Vlasenko and especially A. Okounkov for very inspiring discussions. I thank A. Varchenko for introducing me to the subject. This note is the extended version of the talks [Sm23] and [Sm24], I thank the organizers for these opportunities. This work is supported by NSF grants DMS - 2054527 and DMS-2401380.
Appendix A $p$-adic $q$-deformations
In this section we recall standard notations of $q$-calculus. We also obtain results about $p$-adic $q$-gamma function which we we need in this paper.
A.1. $q$-numbers
For $n\in\mathbb{N}$ we define the $q$-number by:
$[n]_{q}=\dfrac{1-q^{n}}{1-q}=1+q+\dots+q^{n-1}.$
Let $[n]_{q}!=[1]_{q}[2]_{q}\dots[n]_{q}$ denote the $q$-factorial and
(A.1) $\binom{n}{k}_{\!\!q}=\dfrac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}$
Let us define the reciprocal of the $q$-gamma functions as formal product:
$(x,q)_{\infty}=\prod\limits_{i=0}^{\infty}(1-xq^{i})$
We also define finite products
$(x,q)_{d}=\dfrac{(x,q)_{\infty}}{(xq^{d},q)_{\infty}}=(1-x)(1-xq)\dots(1-xq^{d-1})$
and
(A.2) $[x,q]_{d}=(1-qx)\dots(1-q^{d-1}x)=\dfrac{(xq,q)_{\infty}}{(xq^{d},q)_{\infty}}$
A.2. $q$-hypergeometric series at unit roots
Let us recall the $q$-binomial theorem:
$\dfrac{(z\hbar,q)_{\infty}}{(z,q)_{\infty}}=\sum\limits_{d=0}^{\infty}\,\dfrac{(\hbar,q)_{d}}{(q,q)_{d}}z^{d}$
At $\hbar=q^{l}$, $l\in\mathbb{N}$ it is convenient to write the coefficients as
(A.3) $\dfrac{(\hbar,q)_{d}}{(q,q)_{d}}=\dfrac{\frac{(\hbar,q)_{\infty}}{(\hbar q^{d},q)_{\infty}}}{\frac{(q,q)_{\infty}}{(qq^{d},q)_{\infty}}}=\dfrac{(\hbar,q)_{\infty}}{(q,q)_{\infty}}\dfrac{(qq^{d},q)_{\infty}}{(\hbar q^{d},q)_{\infty}}$
Using notation (A.2) at $\hbar=q^{l}$ we thus write:
(A.4) $\dfrac{(q^{l},q)_{d}}{(q,q)_{d}}=\dfrac{[q^{d},q]_{l}}{[1,q]_{l}}$
At the same time at $\hbar=q^{l}$ we have
$\dfrac{(z\hbar,q)_{\infty}}{(z,q)_{\infty}}=\dfrac{1}{(z,q)_{l}}$
Combining all this together, we arrive at the following version of the $q$-binomial theorem:
(A.5) $\dfrac{1}{(z,q)_{l}}=\sum\limits_{d=0}^{\infty}\,\dfrac{[q^{d},q]_{l}}{[1,q]_{l}}z^{d}$
Note that the coefficients of the power series (A.5) are polynomials in $q$ and in particular, can be specialized at unit roots. Let $p$ be a prime and let $\zeta\neq 1$ be a unit root $\zeta^{p^{s}}=1$ for some natural number $s$. We may assume that the order of $\zeta$ is $p^{m}$ with $1\leqslant m\leqslant s$. Let $h^{\prime}$ be a natural number. Then, applying (A.5) for $l=h^{\prime}p^{s}$ we have
$\left.\dfrac{1}{(z,q)_{h^{\prime}p^{s}}}\right|_{q=\zeta}=\dfrac{1}{(1-z^{p^{m}})^{h^{\prime}p^{s-m}}}=\sum\limits_{i=0}^{\infty}\,\left.\dfrac{[q^{i},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right|_{q=\zeta}z^{i}$
At the same time, since $\zeta^{p}$ is a primitive unit roots of order $p^{m-1}$ applying (A.5) for $l=h^{\prime}p^{s-1}$ we find
$\left.\dfrac{1}{(z^{p},q^{p})_{h^{\prime}p^{s-1}}}\right|_{q=\zeta}=\dfrac{1}{(1-z^{p^{m}})^{h^{\prime}p^{s-m}}}=\sum\limits_{i=0}^{\infty}\,\left.\dfrac{[q^{pi},q^{p}]_{h^{\prime}p^{s-1}}}{[1,q^{p}]_{h^{\prime}p^{s-1}}}\right|_{q=\zeta}z^{i}$
Comparing the coefficients we conclude
Lemma A.1.
Let $\zeta\neq 1$ be a unit $\zeta^{p^{s}}=1$ and let $h^{\prime}\in\mathbb{N}$ then for any $i\in\mathbb{N}$ we have
(A.6) $\left.\dfrac{[q^{i},q]_{h^{\prime}p^{s}}}{[1,q]_{h^{\prime}p^{s}}}\right|_{q=\zeta}=\left.\dfrac{[q^{pi},q^{p}]_{h^{\prime}p^{s-1}}}{[1,q^{p}]_{h^{\prime}p^{s-1}}}\right|_{q=\zeta}$
Coefficients (A.6) are equal to 00 unless $p\mid i$.
Lemma A.2.
a) If $h\in\mathbb{Z}$ then the hypergeometric function $F(h,\boldsymbol{a},q,z)$ (2.3) does not have poles at $q$ given by roots of unity.
b) Let $k\in\mathbb{Z}$ and $\zeta$ be a primitive unit root of order $m$ then:
$\left.F(mk,\boldsymbol{a},q,z)\right|_{q=\zeta}=\dfrac{1}{(1-z^{m})^{k}}$
Proof.
The proof follows directly from the definition of $q$-hypergeometric series. ∎
A.3. $p$-adic $q$-deformations
We recall the original definitions of the $p$-adic gamma and $q$-gamma functions:
Theorem A.3 (Koblitz, [Ko80]).
For $q=1+t$ with $|t|_{p}<1$ the function defined on $n\in\mathbb{N}$ by $\Gamma_{p,q}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,[i]_{q}$ extends to a continuous function on $\mathbb{Z}_{p}$.
In the limit $t\to 0$ we have $\lim\limits_{q\to 1}\,\Gamma_{p,q}=\Gamma_{p}$ where $\Gamma_{p}$ denotes the $p$-adic gamma function of Morita [Mo75], which is the extension to $\mathbb{Z}_{p}$ of $\Gamma_{p}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,n.$
Lemma A.4.
For $q=1+t$ with $|t|_{p}<1$, the function of $n,h\in\mathbb{N}$ given by
(A.7) $f(h,n)=\dfrac{\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}}{\dfrac{(q^{ph},q^{p})_{\infty}}{(q^{p},q^{p})_{\infty}}\dfrac{(q^{pn+p},q^{p})_{\infty}}{(q^{pn+ph},q^{p})_{\infty}}}$
extends to $\mathbb{Z}_{p}$ values of the arguments by $f(h,n)=\dfrac{\Gamma_{p,q}(pn+ph)}{\Gamma_{p,q}(pn)\,\Gamma_{p,q}(ph)}$.
Proof.
If $n$ and $h$ are two natural numbers then
$\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}=\binom{pn+ph-1}{pn}_{\!\!q}$
is a $q$-binomial coefficient (A.1). Similarly
For the ratio of these factors we find:
By Theorem A.3 this function extends to $\mathbb{Z}_{p}$-values of $h$ and $n$. ∎
Let us recall that the $p$-adic unit roots $q\in\Omega$ of order $p^{s}$, are of the form $q=1+t$ with $|t|_{p}=p^{-\frac{1}{p^{s-1}(p-1)}}<1$, [Ko84]. By Theorem A.3 we can consider the values of $p$-adic $q$-gamma functions at such $q$. The following Lemma will be useful in the next section.
Lemma A.5.
let $x,y\in\mathbb{Z}_{p}$ and $q\neq 1$ be a $p^{s}$-th unit root, then
(A.8) $\Gamma_{p,q}(x+yp^{s})=\Gamma_{p,q}(x)\Gamma_{p,q}(yp^{s})$
Proof.
First we note that for $x\in\mathbb{Z}_{p}$ and $q\neq 1$ given by a $p^{s}$-th unit root we have:
(A.9) $\Gamma_{p,q}(x+p^{s})=\Gamma_{p,q}(x)\Gamma_{p,q}(p^{s})$
Indeed, for $x\in\mathbb{N}$ the identify (A.9) follows directly from the definition of $(p,q)$-Gamma function:
$\Gamma_{p,q}(x)=(-1)^{n}\prod\limits_{{0<i<x}\atop{p\nmid i}}\,[i]_{q}$
which extends to $x\in\mathbb{Z}_{p}$ as in Theorem A.3. Next, for $y\in\mathbb{N}$ (A.8) follows by applying (A.9) $y$-times. Then both sides of (A.8) extend to $y\in\mathbb{Z}_{p}$ as in Theorem A.3 again. ∎
A.4.
We can write
$\Gamma_{p,q}(p)=-\dfrac{1-q}{1-q}\dfrac{1-q^{2}}{1-q}\cdots\dfrac{1-q^{p-1}}{1-q}=-\lim\limits_{w\rightarrow 1}\,\dfrac{1-w}{1-w}\dfrac{1-wq}{1-q}\cdots\dfrac{1-wq^{p-1}}{1-q}$
where the limit is for regularization purposes. Now let us assume that $q$ is a primitive $p$-th root of unity, $q^{p}=1$ then from above formula we obtain:
$\Gamma_{p,q}(p)=\dfrac{-1}{(1-q)^{p-1}}\lim\limits_{w\rightarrow 1}\,\dfrac{1-w^{p}}{1-w}=\dfrac{-p}{(1-q)^{p-1}}$
Using same logic we also find for $m\in\mathbb{N}$:
$\Gamma_{p,q}(mp)=\Gamma_{p,q}(p)^{m}=\Big{(}\dfrac{-p}{(1-q)^{p-1}}\Big{)}^{m}$
For primitive $p^{s}$-th we can use a similar trick. First we write:
$\Gamma_{p,q}(p^{s})=-\prod\limits_{{0<i<p^{s}}\atop{p\nmid i}}\,\dfrac{1-q^{i}}{1-q}=\dfrac{-1}{(1-q)^{p^{s-1}(p-1)}}\lim\limits_{w\to 1}\,\dfrac{\prod\limits_{{0<i<p^{s}}}(1-wq^{i})}{\prod\limits_{{0<i<p^{s-1}}}(1-wq^{ip})}$
Assume $q$ is a primitive $p^{s}$-th root of unity, $q^{p^{s}}=1$, then the previous expression gives:
$\Gamma_{p,q}(p^{s})=\dfrac{-1}{(1-q)^{p^{s-1}(p-1)}}\lim\limits_{w\to 1}\dfrac{1-w^{p^{s}}}{1-w^{p^{s-1}}}=\dfrac{-p}{(1-q)^{p^{s-1}(p-1)}}$
In general we conclude
Proposition A.6.
Let $q$ be a primitive $p^{s}$-th unit root and $m\in\mathbb{N}$ then
$\Gamma_{p,q}(mp^{s})=\Big{(}\dfrac{-p}{(1-q)^{p^{s-1}(p-1)}}\Big{)}^{m}$
Note also that a primitive $p^{s}$-th unit root has the form $q=1+t$ with $|t|_{p}=p^{-\frac{1}{p^{s-1}(p-1)}}$. Thus we see that $|\Gamma_{p,q}(p^{s})|_{p}=1$.